A glass sphere with radius 4.00 mm, mass 55.0 g, and total charge 2.00 µC is separated by 150.0 cm from a second glass sphere 2.00 mm in radius, with mass 275.0 g and total charge −5.00 µC. The charge distribution on both spheres is uniform. If the spheres are released from rest, what is the speed of each sphere the instant before they collide?

v1=?

v2=?
Answer #1

a)

By Conservation of momentum

m1v1-m2v2=0

=>v2=(m1/m2)v1

By Conservation of energy

-Kq1q2/d =(1/2)m1v12+(1/2)m2v22-Kq1q2/(r1+r2)

(9*109)(2*10-6)(5*10-6)[-1/1.5+1/(0.004+0.002)] =(1/2)(0.055)v12+(1/2)(0.275)(0.055/0.275)2v12

v1=21.28 m/s

b)

v2=(55/275)(10.824) =4.255 m/s